OpenStudy (anonymous):

write the mixed number as an inproper fraction 5 3/4 1 1/12

OpenStudy (anonymous):

5 3/4 = 23/4 1 1/12 = 13/12

OpenStudy (anonymous):

lol221 and i have the same answer because 5 3/4 =5*4 +3=23 the copy the denominator 1 1/12=1*12+1=13 then copy the denominator